# Controllable Motion Generation via Diffusion Modal Coupling

### v1

@@ Abstract
Diffusion models have recently gained significant attention in robotics due to their ability to generate multi-modal distributions of system states and behaviors. However, a key challenge remains: ensuring precise control over the generated outcomes without compromising realism. This is crucial for applications such as motion planning or trajectory forecasting, where adherence to physical constraints and task-specific objectives is essential. We propose a novel framework that enhances controllability in diffusion models by leveraging multi-modal prior distributions and enforcing strong modal coupling. This allows us to initiate the denoising process directly from distinct prior modes that correspond to different possible system behaviors, ensuring sampling to align with the training distribution. We evaluate our approach on motion prediction using the Waymo dataset and multi-task control in Maze2D environments. Experimental results show that our framework outperforms both guidance-based techniques and conditioned models with unimodal priors, achieving superior fidelity, diversity, and controllability, even in the absence of explicit conditioning. Overall, our approach provides a more reliable and scalable solution for controllable motion generation in robotics.
@@ I INTRODUCTION
Diffusion models [1] have recently emerged as a powerful class of deep generative models, combining stability, multi-modal expressiveness, and high-fidelity sample generation. These attributes make them well-suited to address the complexity and uncertainty prevalent in robotics applications. Indeed, their impact spans diverse domains, ranging from sensor simulation [2], e.g. synthesizing realistic camera data for domain randomization and sim-to-real transfer [3]), to trajectory generation [4] where they produce diverse motion plans under challenging environmental conditions [5]), and policy learning [6], wherein they model multi-step action distributions for high-dimensional control tasks [7]). These applications highlight how diffusion-based approaches can reshape core robotics paradigms, facilitating robust and adaptive robotic systems through high-fidelity probabilistic modeling.
Figure 1: High-level comparison between guidance-based and our proposed methods. (a) A standard diffusion model captures a multi-modal data distribution. For illustrative purposes, only three modes are visualized. A guidance function (in red) is employed to direct the generation towards mode II, i.e., the less likely yet operational behaviors. (b) When the guidance factor is low, it fails to effectively steer the model, which continues to produce the more probable, dominant modes. (c) Conversely, when the guidance factor is high, sample fidelity and feasibility degrades. Because guidance is applied to intermediate predictions during denoising, it can push these predictions out of the data manifold’s high-fidelity region, causing distribution match. (d)(e) In our method, the less likely data mode is tightly coupled to a corresponding prior mode. During sampling, we simply initiate the denoising process from that specific prior mode, achieving direct controllability without sacrificing sample fidelity.
However, several challenges persist, and the chief one among them is controllability. Generating samples that adhere to specific objectives and domain constraints is nontrivial. Naive sampling may yield physically implausible trajectories or policies lacking necessary structure. Even worse, when controllability is insufficient, practitioners often generate large numbers of samples to ensure they capture desirable outcomes. This poses scalability concerns and adds challenges to mine high-fidelity samples for large-scale synthetic generations.
Existing solutions commonly resort to constraint-based sampling or post-hoc guidance [8, 9]. The core idea is to incorporate domain knowledge and task-specific constraints during the sampling process. While such methods indeed promote better alignment with target objectives, they risk degrading sample fidelity by shifting outputs away from the learned high-fidelity data manifold. This is because manually imposed constraints or guidance signals risk overriding the model’s intrinsic generative dynamics. Consequently, suboptimal or unnatural samples that no longer reflect the true data distribution would be led. Balancing fine-grained control with the preservation of sample fidelity thus remains an open research problem (Figure 1).
@@
Figure 2: High-level overview. Conventional diffusion models rely on a unimodal prior distribution, providing no inherent mechanism to control which trajectories get emphasized. This limitation can neglect lower-probability yet operationally crucial motion plans - an important limitation for robotics applications that demand flexible behaviors. In our work, we propose a multi-modal prior distribution and enforcing strong modal coupling between prior and data distributions through a novel diffusion process. By denoising directly from the relevant prior mode, we preserve trajectory fidelity and realism without risking distribution mismatch. Crucially, even with an unconditioned diffusion model, our framework offers straightforward control over which data mode to attend, facilitating precise and adaptive motion generation. In the figure, each prior corresponds to one system behavior. “ACC”, “DEC” and “MSP” refer to speed modes (acceleration, deceleration, and maintaining speed), while “R”, “L” and “S” represent steering modes (right, left, and straight).
In this work, we propose a novel framework that achieves strong controllability in diffusion models, without incurring the distribution drift commonly encountered in guidance-based approaches. The key idea is to replace the unimodal prior in a standard diffusion model with a multi-modal distribution that is tightly coupled to the principle modes of the target data. Therefore at sampling, rather than applying post-hoc guidance to steer the generative process toward a specific mode, we initiate the denoising process directly from the corresponding prior mode. This ensures that sampling remains inherently aligned with the training distribution, eliminating the need for external guidance and mitigating the risk of distribution mismatch (Figure 2).
@@
Our current formulation assumes that the data modes subject to control are explicitly known, yet the diffusion model itself does not require conditioning on these mode labels. While this assumption may appear restrictive, it serves as a foundation step towards enabling strong controllability over data with unknown key modes, wherein the central challenge shifts to accurately identifying the appropriate prior mode based on target constraints. Nevertheless, we demonstrate that by adopting a multi-modal prior distribution, strong modal coupling, and a careful prior parametrization, our method significantly outperforms guidance-based techniques and even conditioned modeling with a unimodal prior, in terms of both fidelity and controllability. We validate these claims on the Waymo dataset for motion prediction, and in Maze2D for multi-task control. The paper is organized as follows. Section II and III cover the related work and the necessary background. We detail the proposed method in Section IV. Section V presents the experimental results, and Section VI concludes the paper.
@@ II Related Work
The typical strategy to control diffusion-based generation is incorporating domain-specific objectives into the generation process. Classifier guidance [13] guides the diffusion model with a separate cost function that encodes the objectives during sampling. Recent works [4] propose different analytical guidance functions to achieve either realism or safety-critical objectives. On the other hand, Classifier-free guidance [8] additionally optimizes a time-dependent conditional model to obtain guidance. For instance, VBD [12] achieves this using the motion predictor jointly optimized for goal-guided trajectory generation. However, the constraints imposed by guidance often degrade the realism of model generation due to distribution mismatch. Our approach completely avoids this issue by coupling prior and data modes.
@@ III Background
Diffusion models are probabilistic generative models that synthesize new data by iteratively denoising an initial noise sample. These models begin by defining a forward stochastic process that progressively adds noise to real data $x_{0}$, eventually converting it into pure noise. Formally, this forward noising process is defined as:
@@
where $x_{1:T}$ denotes the sequence of noised samples from step $1$ to $T$, $\epsilon_{t}$ is standard Gaussian noise, and $\beta_{t}$ is the pre-defined forward variance. Let $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$. This forward process allows one to directly sample $x_{t}$ at any intermediate step $t$ in close form:
@@
In a typical diffusion setup, the variance schedule $\beta_{1:T}$ is chosen such that $\bar{\alpha}_{T}$ approaches 00 under a sufficiently large $T$. This ensures the final noised state $x_{T}$ to converge to pure noise, establishing a standard unimodal Gaussian prior.
A reverse process recovers clean samples by removing noise step by step, modeled as a Gaussian distribution whose mean is given by a neural network and whose variance is fixed based on the forward variance schedule. Concretely,
@@
where $\theta$ denotes the parameters of the neural denoiser trained to maximize the variational lower bound [1] on the log-likelihood of the observed data $x_{0}$: $\max_{\theta}$ $-\log p_{\theta}(x_{0}|x_{1})$ $+$ $\sum_{t}D_{KL}[q(x_{t-1}|x_{t},x_{0})||p_{\theta}(x_{t-1}|x_{t})]$. To simplify learning, one can reparameterize $\mu_{\theta}$ in terms of noise prediction: $\mu_{\theta}(x_{t})=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$, where $\epsilon_{\theta}$ is optimized to predict the true noise, approximating the score, i.e., the gradient of log density $\nabla_{x_{t}}\log q(x_{t})$, across different noise scales. Alternatively, one can predict the clean sample $x_{\theta}^{0}$, using it directly to derive close-form posterior mean:
@@
Controllability in diffusion models refers to the ability to impose particular constraints on the generated samples. A common strategy for achieving this is through guidance [8], where an additional cost function $f$ is integrated into each step of the reverse denoising process. Formally, the learned scores are modified as follows:
@@
where $\omega$ controls the influence of $f$. This cost function may encode a variety of task-specific constraints, such as enforcing speed limits or steering preferences in autonomous driving, restricting actions in control tasks, or nudging the model towards a particular label or style in image generation.
@@ IV Enhanced Diffusion Controllability via Modal Coupling
A common approach to controlling diffusion-based generation is through guidance. However, as the guidance function is not integrated into the core training objective, it imposes post-hoc constraints that may incur a distribution mismatch, causing samples to drift away from high-fidelity regions of the data manifold.
To address these challenges, we propose adopting a multi-modal prior, and pose the following question: if each prior mode is coupled with a corresponding mode in the target distribution, can we run the reverse process from different prior modes for strong controllability without guidance? Not only is the answer “yes”, but it also promises to eliminate reliance on post-hoc constraints. As a result, each mode of the target distribution is naturally accommodated during sampling, with no mismatch risk posed by guidance.
Figure 3: 2D toy example. (a) The data distribution with eight distinct modes. (b) Results from DDPM [1], highlighting spurious samples in the gaps between modes. (c)-(d) Results from our method, which trains an unconditioned diffusion model but explicitly pairs each prior mode with a data mode. This yields significantly fewer spurious samples and provides direct control over which mode is generated. (e)-(f) The pre-defined prior parametrizations. In particular, (e) shows that when the prior means lie far from the origin (i.e., have large magnitudes), the diffusion model struggles to learn those corresponding modes effectively.
In this work, we assume that the modes of data distribution are explicitly known, and model the multi-modal prior as a Gaussian mixture model:
@@
where $k$ is the number of modes, $r_{i}$ denotes the proportion of data with mode label $i$, and $\mu_{i}$ and $\sigma_{i}^{2}$ specify the mean and variance of the $i$-th Gaussian component, respectively. Notably, when conditioning on a specific data $x_{0}$ with label $L$, the prior reduces to a unimodal Gaussian:
$x_{T}|x_{0}\sim\mathcal{N}(\mu_{L},\sigma_{L}^{2}I).$ (9)
This structure allows us to explicitly account for different modes in the data while retaining a unimodal form given a specific label. In the sequel, we will demonstrate that, by coupling these prior modes with the corresponding data modes and by carefully parameterizing each prior mode, we can achieve robust mode control even using unconditioned diffusion models.
IV-A Modal Coupling
We begin by defining forward and reverse diffusion processes that accommodate a general Gaussian prior $x_{T}|x_{0}\sim\mathcal{N}(\mu,\sigma^{2}I)$.
@@ IV-A Modal Coupling
The proof is in Appendix VI-A. Lemma 1 establishes that introducing the constant shift term $\mu/\eta_{T}$ at each forward step and scaling the Gaussian noise by $\sigma$ ensure the final marginal $q(x_{T}|x_{0})$ to align with the desired Gaussian prior, which sets the foundation for the corresponding reverse process.
@@
We include the proof in Appendix VI-B. This concludes the derivation of the proposed forward and reverse processes that support an arbitrary Gaussian, thereby enabling any prior mode to be directly coupled with a specific mode in the target distribution. Given a labeled dataset $\mathcal{X}$ and a set of pre-defined prior mode parameterizations, we train the diffusion model $x_{\theta}^{0}$ by:
@@
Here, we adopt the reparameterization of $\mu_{\theta}$, i.e., the mean of posterior $p(x_{t-1}|x_{t})$, in term of clean sample prediction. Switching to noise prediction is straightforward but omitted due to performance considerations.
At sampling, we can either draw from the multi-modal prior distribution (8) to generate across all data modes, or select a unimodal prior (9) to concentrate on a specific mode, allowing fine-grained control over the diffusion process.

### v2

@@ Abstract
Diffusion models are increasingly used in robotics to represent multi-modal distributions over system states and behaviors, but precise control of generated outcomes without degrading physical realism remains challenging. This paper introduces a controllable diffusion framework that (i) replaces the standard unimodal Gaussian prior with an explicit multi-modal prior, and (ii) enforces modal coupling between prior components and principal data modes through novel forward and reverse diffusion processes. Sampling is initialized directly from a selected prior mode aligned with task constraints, avoiding train–test mismatch and manifold drift commonly induced by post‑hoc guidance. Empirical evaluations on motion prediction (Waymo Dataset) and multi‑task control (Maze2D) show consistent improvements over guidance‑based baselines in fidelity, diversity, and controllability. These results indicate that multi‑modal priors with strong modal coupling provide a scalable basis for controllable motion generation in robotics.
@@ I INTRODUCTION
Diffusion models [1] provide expressive, high-fidelity generative capabilities well suited to robotics, where uncertainty and multi-modality are intrinsic. Recent applications span sensor simulation [2, 3], trajectory generation [4, 5], and control policy learning [6, 7], illustrating the utility of diffusion for modeling complex, high-dimensional behavior distributions. Despite this progress, controllability, i.e., the ability to produce samples that satisfy task objectives and domain constraints while remaining on the data manifold, remains a central obstacle. Unconstrained sampling can produce implausible plans or policies, thereby undermining the intended driving objectives [4]. Due to limited controllability, practitioners resort to generating large numbers of samples to capture desirable outcomes, which raises scalability concerns and complicates the mining of high-fidelity samples for large-scale synthetic generations.
Figure 1: High-level comparison of guidance-based approaches versus our proposed method. (a) A standard diffusion model fits a multi-modal data distribution (three modes). A guidance term (red) attempts to steer sampling toward a rare yet operationally critical mode (Mode II). (b) Standard (unguided) sampling concentrates on high-probability modes. (c) Guidance perturbs intermediate states off the well-trained data manifold, degrading fidelity. (d)(e) Our method couples each principal mode to a dedicated prior component, enabling direct, mode-aligned control at sampling while avoiding guidance-induced distribution mismatch.
Existing approaches typically inject constraints at inference, via constraint-based sampling or post-hoc guidance, to encode domain knowledge and task objectives into the reverse process [8, 9],. While such mechanisms improve alignment with target objectives, they introduce a train-test mismatch: external gradients modify the learned score field during sampling, perturbing the reverse dynamics and steering intermidiate states away from high‑density regions of the data distribution. The resulting off-manifold drift degrades sample fidelity. Achieving fine-grained controllability without sacrificing data-manifold realism therefore remains an open problem (Figure 1 [a-c]).
@@
Figure 2: High-level overview. Conventional diffusion models use a unimodal prior distribution and lack an intrinsic mechanism to select which trajectories are emphasized. We introduce a multi-modal prior and enforce strong modal coupling between prior and data via a novel diffusion process. The framework enables direct mode selection even with an unconditioned diffusion model, supporting precise and adaptive motion generation. In the figure, each prior component corresponds to one behavior. “ACC”, “DEC” and “MSP” refer to speed modes (acceleration, deceleration, and maintaining speed), while “R”, “L” and “S” represent steering modes (right, left, and straight).
We address this problem by introducing modal coupling via a multi-modal prior. Concretely, we replace the standard Gaussian prior with a Gaussian-mixture prior whose components are in one-to-one correspondence with principal modes of the data distribution (Figure 1 [d-e]). We derive modified forward and reverse diffusion processes that (i) map all data samples sharing the same label to a common, non-standard Gaussian during noising and (ii) recover the corresponding data mode when denoising is initialized from the associated prior component. At inference, controllability is obtained by selecting the prior mode consistent with the target constraints and running standard denoising, without post-hoc guidance. By removing external guidance terms, our method eliminates train–test mismatch and mitigates intermediate-step distribution drift, while retaining flexible controllability (Fig. 2).
@@
Our current formulation assumes that the data modes subject to control are explicitly known, yet the diffusion model itself does not require conditioning on these mode labels. While this may appear restrictive, it provides a foundation for strong controllability over data with unknown key modes, wherein the central challenge shifts to accurately identifying the appropriate prior mode from target constraints. Nevertheless, by adopting a multi-modal prior distribution, strong modal coupling, and a carefully-designed prior parametrization, our method significantly outperforms guidance-based techniques in both fidelity and controllability. We validate these claims on the Waymo dataset for motion prediction, and in Maze2D for multi-task control. The paper is organized as follows. Section II and III review related work and background. We detail the proposed method in Section IV. Section V presents the experimental results, and Section VI concludes.
@@ II Related Work
The typical strategy to control diffusion-based generation is incorporating domain-specific objectives into the generation process. Classifier guidance [13] guides the diffusion model with a separate cost function that encodes the objectives during sampling. Recent works [4, 14, 15, 16] introduce either language-model–generated or analytical guidance functions to pursue realism or safety-critical objectives.
On the other hand, Classifier-free guidance [8] additionally optimizes a time-dependent conditional model to obtain guidance. This is widely adopted in other fields including text-to-image [17] and 3D objection [18] generation. However, the constraints imposed by guidance often degrade the realism of generated motion [14] due to distribution mismatch between denoising steps and corresponding diffusion forward steps. Our approach avoids this issue by coupling prior and data modes and constructing a shared probability path for both forward and reverse diffusion.
@@ III Background
Diffusion models are probabilistic generative models that synthesize new data by iteratively denoising an initial noise sample. They first define a forward stochastic process that progressively perturbs real data $x_{0}$ into pure noise:
@@
where $x_{1:T}$ denotes the sequence of noised samples, $\epsilon_{t}$ is standard Gaussian noise, and $\beta_{t}$ is forward variance. Let $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$. The above forward process yields the closed-form marginal:
@@
Choosing a variance schedule $\beta_{1:T}$ with $\bar{\alpha}_{T}\to 0$ ensures the prior $x_{T}\sim\mathcal{N}(0,I)$, i.e., a standard Gaussian distribution.
The reverse process removes noise step by step via Gaussian transitions. Concretely,
@@
where $\theta$ are the denoiser parameters trained by maximizing the evidence lower bound [1]: $\max_{\theta}$ $\log p_{\theta}(x_{0}|x_{1})$ ++ $\sum_{t}D_{KL}[q(x_{t-1}|x_{t},x_{0})||p_{\theta}(x_{t-1}|x_{t})]$. A common reparameterization predicts noise:
$\mu_{\theta}(x_{t})=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)),$ (6)
where $\epsilon_{\theta}$ approximates the score $\nabla_{x_{t}}\log q(x_{t})$ across noise scales. Alternatively, predicting the clean sample $x_{\theta}^{0}$ gives the closed-form posterior mean:
@@
Controllability denotes the ability to enforce the specified constraints on generated samples. A common approach is guidance [8] which augments the learned score at each reverse step with a task-specific cost $f$:
@@
where $\omega$ controls the strength of the guidance. Here, $f$ is flexible: it can encode hand-designed differentiable penalties (e.g., speed limits, action bounds, or label/style preferences) or be realized implicitly via a conditional score model, as in the popular classifier-free guidance approaches [19].
@@ IV Enhanced Diffusion Controllability via Modal Coupling
A common strategy for controlling diffusion‑based generation is guidance. However, guidance is absent from diffusion model training and is applied only at sampling as a post-hoc modification (8). This train–test mismatch can induce distribution shift, pushing intermediate states off the high‑fidelity data manifold. To address this, we propose adopting a multi-modal prior and pose the question:
if each prior mode is tightly coupled to a principal data mode at training, can we run denoising process from different prior modes for direct controllability without guidance?
We show the answer is yes: by modifying the diffusion processes to accommodate a carefully-parameterized Gaussian-mixture prior, we can initialize the denoising/reverse process from a selected prior mode to target specific data constraints. Importantly, this eliminates reliance on post-hoc modifications, avoiding guidance-induced distribution mismatch.
Figure 3: 2D toy example. (a) The data distribution with four distinct modes. (b-c) Results from DDPM [1] show using unimodal prior yields spurious samples in the gaps between modes. (d-e) When prior means have large magnitude (i.e., lie far from the origin), the diffusion model struggles to recover realistic per-mode data distributions. (f-g) Insufficient separation between prior modes also prevents the model from accurately capturing the data distribution. (h-i) With a carefully designed prior parameterization that maintains clear separation between modes without introducing excessive values, our method produces substantially fewer spurious samples and further enables direct control over individual modes. Corresponding modes and samples share the same color.
IV-A Gaussian Mixture as a Multi-Model Prior
We assume data modes are explicitly known, and model the multi-modal prior as a Gaussian mixture model:
@@ IV-A Gaussian Mixture as a Multi-Model Prior
where $k$ is the number of modes, $r_{i}$ is the proportion of data with mode label $i$, and $\mu_{i}$, $\sigma_{i}^{2}$ are the mean and variance of the component $i$. Conditioned on a specific label $L$, the prior reduces to a unimodal Gaussian:
$x_{T}|x_{0}\sim\mathcal{N}(\mu_{L},\sigma_{L}^{2}I),$ (10)
allowing us to explicitly account for different data modes while retaining a unimodal form given a specific label.
IV-B Modal Coupling
Effective training with a Gaussian mixture prior requires two conditions. First (modal coupling): each prior mode must correspond one-to-one with a data mode. For label $L$, the forward process maps data with label $L$ to $\mathcal{N}(\mu_{L},\sigma_{L}^{2}I)$, while initializing the reverse process from that specific prior mode recovers the data of label $L$. Second (trajectory separation): throughout denoising, diffusion paths from distinct prior modes must remain sufficiently separated to preserve mode identity and avoid cross-mode interference. This section focuses on modal coupling, while the discussion of prior parameterization for trajectory separation is in Section IV-C.
We first define a forward noising process such that the terminal distribution satisfies a general Gaussian prior $x_{T}|x_{0}\sim\mathcal{N}(\mu,\sigma^{2}I)$.
@@ IV-B Modal Coupling
Proof. See Appendix VI-A. It shows that a constant shift term $\mu/\eta_{T}$ at each forward step and noise scaling by $\sigma$ yield the desired terminal prior, enabling a matched reverse process.
@@
Proof. See Appendix VI-B. This completes the derivation of the modified forward and reverse processes, making each data mode directly associated with a non-standard Gaussian prior component. We now introduce the training objective. Given a labeled dataset $\mathcal{X}$ and pre-defined prior parameters $\{\mu_{1:k},\sigma_{1:k}\}$, we construct noisy samples via (12) and train the clean-prediction model $x_{\theta}^{0}$ with:
@@
Switching the reparameterization of $\mu_{\theta}$ to noise prediction is straightforward but omitted for performance considerations.
At sampling, we can draw the prior sample $x_{T}$ from the mixture (9) to cover all modes, or fix a component (10) to target specific constraints.
IV-C Prior Parametrization
Ensuring trajectory separation is crucial for reliable mode identification, which in turn underpins controllability. Without it, reverse trajectories from different prior components may overlap, causing mode indistinguishability and loss of control (Figure 3). We address this by carefully parameterizing the Gaussian mixture prior.
To place the Gaussian means, we draw on the concept of placing evenly spaced points on a high-dimensional sphere. Let $d$ be the data dimension, and $k$ the number of modes. If $k\leq d+1$, which is common in complex data distributions specialized by diffusion models, the means can be located at the vertices of a (k1)(k-1)-simplex embedded in $\mathbb{R}^{d}$ [20]. Let $e_{1:k}\subset\mathbb{R}^{k}$ be the standard basis vectors and $\mathds{1}_{k}$ the all‑ones vector. With sphere radius $\delta>0$, define
$w_{i}$ $=\delta\cdot\sqrt{\frac{k}{k+1}}\cdot(e_{i}-\mathds{1}_{k}\cdot\frac{1}{k})$ (18)
$\mu_{i}$ $=[w_{i}^{1},...,w_{i}^{k},0,...,0]^{T}\in\mathds{R}^{d},$ (19)
where $w_{i}^{j}$ denotes the $j$-th component of $w_{i}$. One can verify that $\|\mu_{i}\|^{2}=\delta$ for all $i$, and the pairwise distance $d$ is:
$d=\delta\cdot\sqrt{2+2/(k-1)}.$ (20)
Finally, we choose the Gaussian variances so that $c$-level confidence ellipsoids do not overlap. That is, for all $i\in[1,k]$,
$\sigma_{i}\sqrt{\mathcal{X}^{2}_{d,c}}\leq d,$ (21)
where $\mathcal{X}^{2}_{d,c}$ is the $c$ quantile of the chi‐square distribution with $d$ degress of freedom [21].
V Experiments